Why does this not follow normal rule? 25-t=15
Is this because the varible is not first.
Please and explain the answer which should be 10.
Thanks.

Found 2 solutions by oscargut, MathTherapy:
Answer by oscargut(2103)

25-t=15
25 - 15= t
t = 10
If you need more help you can contact me at
mthman@gmail.com
Thanks

Answer by MathTherapy(10555)


Why does this not follow normal rule? 25-t=15
Is this because the varible is not first.
Please and explain the answer which should be 10.
Thanks.
 What "normal" rule doesn't the solution for this equation follow?